OpenStudy (anonymous):

how many degress of a minute han of a clock move through when the time changes from 8:00 to 8:27..please explain

OpenStudy (amistre64):

60/27 = 360/?

OpenStudy (amistre64):

there are 60 minutes around a clock which equal 360 degree in a circle

OpenStudy (amistre64):

the 27th part of 60 will equal the nth degree of 360

OpenStudy (amistre64):

another way to see it: 60 n 360 -- * -- = --- 27 n x 60*n = 360 n = 360/60 = 6 27 * 6 = x